# Similarity-Distance-Magnitude Activations

### v2

@@ 5 Results
Co-variate-shifted and Out-of-distribution data.
The bottom half of Table 1 provides the results for $\operatorname{\text{{SentimentOOD}}}$. Here, the distinctions between the estimators become clear, with the non-$\operatorname{\text{{sdm}}}$-based estimators performing poorly, even in terms of the marginal accuracy. That would come as a surprise to end-users, whereas with the $\operatorname{\text{{sdm}}}$ and $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ estimators, the out-of-distribution documents are more reliably rejected, with the few admitted predictions generally obtaining high conditional accuracies, despite the relatively low accuracies of the test set without selection (see $\operatorname{\text{{no-reject}}}$). A similar pattern is observed over the $\operatorname{\text{{Factcheck}}}$ dataset in Table 2.
Understanding $q^{\prime}_{\rm{min}}$.
For reference, Table 3 provides the results over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ for the $\operatorname{\text{{sdm}}}$-based estimators. The value of $q^{\prime}_{\rm{min}}$ tends to increase as the accuracy over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ decreases, reflecting a more conservative $\operatorname{\text{{high-reliability}}}$ region that admits fewer points. Alg. 1 failed to find a finite $q^{\prime}_{\rm{min}}$ for $\operatorname{\text{{Mixtral8x7B+sdm}}}$ over the $\operatorname{\text{{Factcheck}}}$ calibration set at $\alpha=0.95$, so for reference, we also show the $\operatorname{\text{{high-reliability}}}$ region at $\alpha=0.94$, as well as in Table 2. In this way, $q^{\prime}_{\rm{min}}$ provides a principled, data-driven indicator of the reliability of the estimates, which is interpretable as a simple indicator as to whether the conditional accuracies are, or are not, obtainable over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ at the specified $\alpha$.
Far out-of-distribution data.
The Appendix provides Table 4 and Table 5 with analogous results for the $\operatorname{\text{{SentimentShuffled}}}$, $\operatorname{\text{{SentimentOODShuffled}}}$, and $\operatorname{\text{{FactcheckShuffled}}}$ datasets. The selection criteria of Eq. 9 reliably rejects the challenging predictions, whereas the non-$\operatorname{\text{{sdm}}}$-based estimators fare poorly, in general. In this way, the $\operatorname{\text{{sdm}}}$ activation serves as an effective out-of-distribution detection method. With existing methods, defining an out-of-distribution point has been task- and problem-specific, and generally challenging over high-dimensional inputs, typically requiring additional modeling beyond that of the calibration or selection method. In contrast, Eq. 9 provides a principled approach for determining such cut-offs in a data- and model-driven manner, with minimal hyper-parameters, resulting in a separation of points over which the estimator tends to be reliable (namely, the admitted points) and those over which the estimates themselves tend to be unreliable.
6 Conclusion
Appendix A Appendix
We provide the results for the far out-of-distribution (OOD) shuffled datasets in Appendix A.1. Additional implementation details are included in Appendix A.2. Appendix A.3 provides an approach for analyzing the effective sample size for both the class-conditional and prediction-conditional estimates.
A.1 Far OOD Shuffled Datasets
Table 4 shows results for the $\operatorname{\text{{SentimentShuffled}}}$ and $\operatorname{\text{{SentimentOODShuffled}}}$ datasets, and Table 5 shows results for the $\operatorname{\text{{FactcheckShuffled}}}$ datasets, as discussed in Section 5.
A.2 Additional Implementation Details
We mean center the input to $g$, the 1-D CNN of the $\operatorname{\text{{sdm}}}$ activation layer and the otherwise identical CNN adaptors of the baseline comparison estimators, via the mean and standard deviation over $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$. In all experiments with adaptor layers, $M=1000$ and we use a mini-batch size of 50. We use the Adam optimizer (Kingma & Ba, 2017) (without weight decay) with a learning rate of $1\times 10^{-5}$ for training.
A.2.1 Implementation of the SDM Activation Function
A.2.2 Implementation of the Empirical CDF Function
A.3 Analyzing the Effective Sample Size
Assumption A.1.
$\mathbf{\hat{n}}=\left[|\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}|^{y_{1}}\cdot{\rm{eCDF}}^{y_{1}}_{\rm{ca}}(q^{\prime}),\ldots,|\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}|^{y_{C}}\cdot{\rm{eCDF}}^{y_{C}}_{\rm{ca}}(q^{\prime})\right]$ (10)
where $|\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}|^{y_{c}}$ is the count of calibration set points with true label $y=c$.
For the prediction-conditional estimates, assuming independent and identically distributed (i.i.d.) data, these sample size estimates can be used to construct a band around the empirical CDFs over $\operatorname{\text{$d_{\rm{nearest}}$}}$ (Eq. 5) using the sharp constant (Massart, 1990) of the distribution-free DKW inequality (Dvoretzky et al., 1956), with $\hat{n}_{\rm{min}}$ taken as the minimum among the estimated sample sizes across classes for the test instance:
$\hat{n}_{\rm{min}}$ $=\min\left[\hat{n}_{1},\ldots,\hat{n}_{C}\right]$ (12)
If $\hat{n}_{\rm{min}}=0$, our convention is to set $\epsilon=1$. We then construct the conservative lower and upper counterparts to the distance quantile of Eq. 5:
@@ A.3 Analyzing the Effective Sample Size
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. 0.01 1. <0.01 1. 0.01 1. <0.01 1. 0.01
@@
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. <0.01 R 0. 1. <0.01 R 0. 1. <0.01
@@
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. <0.01 1. <0.01 1. <0.01 1. <0.01 1. 0.01
@@
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. <0.01 0.83 <0.01 0.88 0.01 1. <0.01 0.93 0.01
@@
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ R 0. - 0. R 0. R 0. R 0.
@@
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. 0.01 - 0. 1. 0.01 R 0. 1. 0.01
@@
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}},\alpha=0.94$ 1. 0.01 - 0. 1. 0.01 R 0. 1. 0.01

### v3

@@ Results.
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 1. 0.01 1. <0.01 1. 0.01 1. <0.01 1. 0.01
@@
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 1. <0.01 R 0. 1. <0.01 R 0. 1. <0.01
@@
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 1. <0.01 1. <0.01 1. <0.01 1. <0.01 1. 0.01
@@
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 1. <0.01 0.83 <0.01 0.88 0.01 1. <0.01 0.93 0.01
@@
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ R 0. - 0. R 0. R 0. R 0.
@@
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 1. 0.01 - 0. 1. 0.01 R 0. 1. 0.01
@@
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}},\alpha=0.94$ 1. 0.01 - 0. 1. 0.01 R 0. 1. 0.01